# Many-body Fu-Kane-Mele index

### v1

@@ 5.2.2 Non-triviality of the index for spin Hall states
We must show that the symmetric SRE state $\omega=\omega_{\mathrm{SH}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ has non-trivial index. We have shown above that the defect state $\psi^{+}$ obtained by the Kato generator is either a Kramers singlet or belongs to a Kramers pair. Now, Lemma 5.6 yields in particular an even unitary $V\in{\mathcal{A}}^{\mathrm{al}}$ such that $\psi^{+}=\omega^{+}\circ\mathrm{Ad}[V]$. By Lemma 4.1 we conclude that $\omega^{+}$ is itself either a Kramers singlet or belongs to a Kramers pair. The claim follows by Definitions 3.15 and 4.11. ∎
@@ 6 Proof of the main theorem
Proof of (1). Triviality of the index for symmetric product states is Lemma 4.9. The claim about the spin Hall states is Proposition 5.7.
@@
Acknowledgements. The work of S.B. and M.R. is supported by NSERC of Canada.
@@ A.3 A continuity result for parity of states
Appendix B Flux insertion for free fermions
In this appendix we will study the spin up component ${\mathscr{P}}^{\uparrow}$ of a spin Hall state ${\mathscr{P}}_{SH}={\mathscr{P}}^{\uparrow}\oplus{\mathscr{P}}^{\downarrow}$. Recall that ${\mathscr{P}}^{\uparrow}$ is a projection on the Hilbert space ${\mathcal{K}}_{1}=l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2})$. We fix an orthonormal basis $\{e_{i}\}_{i=1,2}$ of ${\mathbb{C}}^{2}$ so that $\{\delta_{x}\otimes e_{i}\}_{x,i}$ is an orthonormal basis of ${\mathcal{K}}_{1}$. For any operator ${\mathscr{A}}$ on ${\mathcal{K}}_{1}$ we denote by
${\mathscr{A}}(x,i;y,j):=\langle\delta_{x}\otimes e_{i},{\mathscr{A}}\,\delta_{y}\otimes e_{j}\rangle$ (B.1)
@@ Appendix B Flux insertion for free fermions
Since in this appendix we will be exclusively concerned with the spin up component we write ${\mathscr{P}}={\mathscr{P}}^{\uparrow}$ here and only here, dropping the spin label from the notation.
Throughout this appendix we will use the following notations : $L=\{x\in{\mathbb{Z}}^{2}\,:\,x_{1}\leq 0\}$ is the left half plane and $R={\mathbb{Z}}^{2}\setminus L$ is the right half plane. For any subset $S\subset{\mathbb{Z}}^{2}$ we write $\chi_{S}$ for the indicator function on $S$, and $\Pi_{S}$ for the projection on S, i.e. $\Pi_{S}|\delta_{x}\otimes i\rangle=\delta_{x\in S}|\delta_{x}\otimes i\rangle$. We write $l_{L}$ and $l_{R}$ for the left and right horizontal axes respectively, regarded as subsets of ${\mathbb{R}}^{2}$. For any $S\subset{\mathbb{R}}^{2}$ and any $x\in{\mathbb{Z}}^{2}$ we write $\mathrm{dist}(x,S)=\inf_{y\in S}\|x-y\|_{1}$ for the distance from $x$ to the set $S$.
B.1 Locality estimates for Fermi projections
Recall that the Fermi projection ${\mathscr{P}}_{SH}$ describing the spin Hall state is exponentially local and so
$\left|{\mathscr{P}}(x,i;y,j)\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}$ (B.2)
where $C,\eta>0$ are the constants introduced in (2.13).
Lemma B.1.
@@ B.1 Locality estimates for Fermi projections
$\left|{\mathscr{P}}_{\phi}^{h}(x,i;y,j)\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}\quad\text{and}\quad\left|{\mathscr{K}}_{\phi}^{h}(x,i;y,j)\right|\leq c{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}$ (B.3)
for some constants $c,\xi>0$ and for all $\phi\in{\mathbb{R}}$, all $x=(x_{1},x_{2}),y=(y_{1},y_{2})\in{\mathbb{Z}}^{2}$ and all $i,j\in\{1,2\}$.
@@
The first claim follows immediately from Eq (B.2) and the definition ${\mathscr{P}}_{\phi}^{h}={\mathrm{e}}^{{\mathrm{i}}\phi\Pi_{h}}{\mathscr{P}}{\mathrm{e}}^{-{\mathrm{i}}\phi\Pi_{h}}$.
In order to bound the matrix elements of ${\mathscr{K}}_{\phi}^{h}={\mathrm{i}}[\partial_{\phi}{\mathscr{P}}_{\phi}^{h},{\mathscr{P}}_{\phi}^{h}]$ we first bound the matrix elements of $\partial_{\phi}{\mathscr{P}}_{\phi}^{h}={\mathrm{i}}[\Pi_{h},{\mathscr{P}}_{\phi}^{h}]$. If $x,y\in h$ of $x,y\notin h$ then $(\partial_{\phi}{\mathscr{P}}_{\phi}^{h})(x,i;y,j)=0$. If $x\in h$ and $y\not\in h$ then using the first claim of this lemma we obtain
$\left|(\partial_{\phi}{\mathscr{P}}_{\phi}^{h})(x,i;y,j)\right|=\left|{\mathscr{P}}_{\phi}(x,i;y,j)\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}\leq C{\mathrm{e}}^{-\eta(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}.$ (B.4)
where for the last inequality we used $x\in h$ and $y\not\in h$ so $\|x-y\|_{1}=\left|x_{1}-y_{1}\right|+\left|x_{2}-y_{2}\right|=\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|$. The same result is obtained if $x\not\in h$ and $y\in h$. The final claim of the lemma now follows straightforwardly from the definition ${\mathscr{K}}_{\phi}^{h}={\mathrm{i}}[\partial_{\phi}{\mathscr{P}}_{\phi}^{h},{\mathscr{P}}_{\phi}^{h}]$ and the bounds on the matrix elements of $\partial_{\phi}{\mathscr{P}}_{\phi}^{h}$ and ${\mathscr{P}}_{\phi}^{h}$. ∎
We now turn to locality properties of the defect states ${\mathscr{P}}_{\phi}$ defined in (5.20). It will be useful to consider the following quantities:
$a_{n}(x,y)$ $:=\sup_{\begin{subarray}{c}\vec{\phi}\in[-\pi,\pi]^{n}\\ i,j\in\{1,2\}\end{subarray}}\,\left|({\mathscr{K}}^{h}_{\phi_{1}}\cdots{\mathscr{K}}^{h}_{\phi_{n}})(x,i;y,j)\right|$ (B.5)
$b_{n}(x,y)$ $:=\sup_{\begin{subarray}{c}\vec{\phi}\in[-\pi,\pi]^{n}\\ i,j\in\{1,2\}\end{subarray}}\,\left|({\mathscr{K}}_{\phi_{1}}\cdots{\mathscr{K}}_{\phi_{n}})(x,i;y,j)\right|.$ (B.6)
Lemma B.2.
There are constants $c,\kappa,\xi>0$ such that
$a_{n}(x,y)$ $\leq c^{n}\kappa^{n-1}{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}$ (B.7)
$b_{n}(x,y)$ $\leq c^{n}\kappa^{n-1}{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}\chi_{L}(x)\chi_{L}(y)$ (B.8)
for all $x,y\in{\mathbb{Z}}^{2}$.

### v2

@@ 5.3 The index extends the Fu-Kane-Mele index for $\operatorname{\mathrm{AII}}$ states
We must show that the $\pi$-flux state $\omega^{+}$ obtained from $\omega_{{\mathscr{P}}}$ by quasi-adiabatic flux insertion is a Kramers singlet if $\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})=0$, and belongs to a Kramers pair of $\operatorname{\mathrm{ind}_{\mathrm{FKM}}}({\mathscr{P}})=1$ (see Definitions 3.15 and 4.11). From Proposition 5.2 we known this to be true for the $\pi$-flux state $\psi^{+}=\omega_{{\mathscr{Q}}^{\mathrm{qa}}_{\pi}}$ obtained from $\omega_{{\mathscr{P}}}$ using a free fermion generator. By Lemma 5.3 there is an even unitary $V\in{\mathcal{A}}^{\mathrm{al}}$ such that $\psi^{+}=\omega^{+}\circ\mathrm{Ad}[V]$. We conclude by Lemma 4.1
@@ 6 Proof of the main theorem
Proof of (1). Triviality of the index for symmetric product states is Lemma 4.9. The claim about the $\operatorname{\mathrm{AII}}$ states is Proposition 5.4.
@@
Acknowledgments. The work of S.B. and M.R. is supported by NSERC of Canada.
Data availability. There is no data associated with this work.
@@ A.3 A continuity result for parity of states
Appendix B Proof that translation invariant $\operatorname{\mathrm{AII}}$ states are stably SRE
Lemma B.1.
Proof.
Since isomorphism classes of complex vector bundles over the torus are completely classified by the Chern number, our assumption $\mathrm{Ch}(\hat{\mathscr{P}})=0$ implies that the bundle $E_{\hat{\mathscr{P}}}$ is isomorphic to a trivial bundle. It now follows from [16, Theorem 23.10], and noting that the torus has a good open cover of 7 sets ([37, Theorem 5.3]), that there is a homotopy $\hat{\mathscr{P}}:[0,1]\times{\mathbb{T}}\rightarrow\mathrm{Gr}_{r}({\mathbb{C}}^{n})$ such that $\hat{\mathscr{P}}_{0}=\hat{P}$ and $\hat{\mathscr{P}}_{1}$ is constant. By [42, Theorem 10.22] we can moreover take this homotopy to be smooth. ∎
$({\mathscr{A}}f)(x)=\frac{1}{4\pi^{2}}\int_{{\mathbb{T}}}\,{\mathrm{d}}k\,\mathrm{e}^{{\mathrm{i}}k\cdot x}\hat{\mathscr{A}}(k)\hat{f}(k)$ (B.1)
$|\langle\delta_{x}\otimes u,{\mathscr{A}}\,\delta_{y}\otimes v\rangle|\leq g_{A}(\mathrm{dist}(x,y))$ (B.2)
Lemma B.2.
Let
$\hat{\mathscr{P}}:[0,1]\times{\mathbb{T}}\rightarrow\mathrm{Proj}_{r}({\mathbb{C}}^{n}):(t,k)\mapsto\hat{\mathscr{P}}_{t}(k)$ (B.3)
${\mathrm{i}}\frac{{\mathrm{d}}{\mathscr{V}}_{t}}{{\mathrm{d}}t}={\mathscr{G}}_{t}{\mathscr{V}}_{t},\qquad{\mathscr{V}}_{0}=\operatorname{\mathbbm{1}},\qquad\text{and}\quad{\mathscr{P}}_{t}={\mathscr{V}}_{t}{\mathscr{P}}_{0}{\mathscr{V}}_{t}^{*}.$ (B.4)
The generators ${\mathscr{G}}_{t}$ have super-polynomial off-diagonal decay in the sense of (B.2), where $g\in{\mathcal{F}}$ is independent of $t$.
Proof.
Let
$\hat{\mathscr{G}}_{t}(k)={\mathrm{i}}\left[\frac{\partial}{\partial t}\hat{\mathscr{P}}_{t}(k),\hat{\mathscr{P}}_{t}(k)\right]$ (B.5)
${\mathrm{i}}\frac{\partial}{\partial t}\hat{\mathscr{V}}_{t}(k)=\hat{\mathscr{G}}_{t}(k)\hat{\mathscr{V}}_{t}(k),\qquad\hat{\mathscr{V}}_{0}(k)=\operatorname{\mathbbm{1}}.$ (B.6)
Then
$\hat{\mathscr{P}}_{t}(k)=\hat{\mathscr{V}}_{t}(k)\hat{\mathscr{P}}_{0}(k)\hat{\mathscr{V}}_{t}(k)^{*}.$ (B.7)
Lemma B.3.
Let $[0,1]\ni t\mapsto{\mathscr{G}}_{t}$ be a norm-continuous family of self-adjoint operators on $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{n})$ such that (B.2) holds uniformly in $t$. Let ${\mathscr{V}}_{t},{\mathscr{P}}_{t}$ be the corresponding one-parameter family of unitaries given by (B.4). Let ${\mathcal{A}}$ be the CAR algebra over $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{n})$ and let $\omega_{t}=\omega_{{\mathscr{P}}_{t}}$ be the gauge-invariant quasi-free state on ${\mathcal{A}}$ corresponding to the projection ${\mathscr{P}}_{t}$, see (2.15). Then there is a $U(1)$ symmetry preserving TDI $F$ such that $\omega_{t}=\omega_{0}\circ\alpha_{t}^{F}$ for all $t\in[0,1]$.
Proof.
${\mathscr{G}}_{t}(x,i;y,j):=\langle\delta_{x}\otimes e_{i},\,{\mathscr{G}}_{t}\,\delta_{y}\otimes e_{j}\rangle$ (B.8)
$F_{x}(t)=\frac{1}{2}\sum_{y\in{\mathbb{Z}}^{2}}\sum_{i,j=1}^{n}\,\left({\mathscr{G}}_{t}(x,i;y,j)a^{*}_{x,i}a_{y,j}+\overline{{\mathscr{G}}_{t}(x,i;y,j)}a^{*}_{y,j}a_{x,i}\right)$ (B.9)
$\sum_{x\in{\mathbb{Z}}^{2}}{\mathrm{i}}\left[F_{x}(t),a(f)\right]=a({\mathrm{i}}{\mathscr{G}}_{t}f)$ (B.10)
which implies that
$\alpha^{F}_{t}(a(f))=a({\mathscr{V}}_{t}^{*}f).$ (B.11)
for any $f\in l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{n})$.
With this, we check that
$(\omega_{0}\circ\alpha^{F}_{t})\big{(}a^{*}(f)a(g)\big{)}=\omega_{0}\big{(}a^{*}({\mathscr{V}}_{t}^{*}f)a({\mathscr{V}}_{t}^{*}g)\big{)}=\left\langle{\mathscr{V}}_{t}^{*}g,{\mathscr{P}}_{0}\,{\mathscr{V}}_{t}^{*}f\right\rangle=\left\langle g,{\mathscr{P}}_{t}\,f\right\rangle=\omega_{t}\big{(}a^{*}(g)a(f)\big{)}.$ (B.12)
We are now equipped to prove Proposition 2.5. We briefly recall the setting described in Section 2.5. The single particle Hilbert space is ${\mathcal{K}}_{m}=l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2m})$ equipped with a fermionic time-reversal ${\mathscr{T}}$. We consider quasi-free states $\omega_{{\mathscr{P}}}$ corresponding to projections ${\mathscr{P}}$ that are exponentially local (2.16), time-reversal invariant (${\mathscr{T}}{\mathscr{P}}{\mathscr{T}}^{*}={\mathscr{P}}$), and translation invariant. It follows that the Fourier transform of ${\mathscr{P}}$ is a smooth projection-valued map $\hat{\mathscr{P}}:{\mathbb{T}}\rightarrow\mathrm{Proj}_{2r}({\mathbb{C}}^{2m})$ with $\mathrm{Ch}(\hat{\mathscr{P}})=0$ for some $0\leq r\leq m$ (recall that the rank of a time-reversal invariant bundle is always even and has vanishing Chern number).
Proof of Proposition 2.5.
If $2m<14r$ then in order to apply Lemma B.1, we expand the ‘ambient dimension’ from $2m$ to $14r$ by considering the algebra ${\mathcal{A}}^{\prime}$ over $l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{2m})\oplus l^{2}({\mathbb{Z}}^{2};{\mathbb{C}}^{14r-2m})$ and the symmetric state $\omega^{\prime}=\omega_{{\mathscr{P}}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ corresponding to the projection ${\mathscr{Q}}={\mathscr{P}}\oplus 0$ and consequently $\hat{\mathscr{Q}}=\hat{\mathscr{P}}\oplus 0$. If $2m>14r$ we simply take ${\mathscr{Q}}={\mathscr{P}}$. The projection-valued map $\hat{\mathscr{Q}}=\hat{\mathscr{P}}\oplus 0$ has $\mathrm{Ch}(\hat{\mathscr{Q}})=0$ and the ambient dimension is large enough so that Lemma B.1 provides a smooth homotopy $(t,k)\mapsto\hat{\mathscr{Q}}_{t}(k)$ interpolating between $\hat{\mathscr{Q}}_{0}=\hat{\mathscr{Q}}$ and a constant projection ${\mathscr{Q}}_{1}$.
Let ${\mathscr{Q}}_{t}$ be the inverse Fourier transform of $\hat{\mathscr{Q}}_{t}$. We then apply Lemma B.2 to obtain unitaries ${\mathscr{V}}_{t}$ and generators ${\mathscr{G}}_{t}$ satisfying the assumption of Lemma B.3. The latter lemma now provides a $U(1)$-invariant TDI $F$ such that $\omega_{t}^{\prime}=\omega_{0}\circ\alpha_{t}^{F}$ where $\omega_{t}^{\prime}=\omega^{\prime}_{{\mathscr{Q}}_{t}}$ is the quasi-free state on ${\mathcal{A}}^{\prime}$ determined by the projection ${\mathscr{Q}}_{t}$. This family of states interpolates between $\omega_{{\mathscr{P}}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ and $\omega_{{\mathscr{Q}}_{1}}$. Since $\hat{\mathscr{Q}}_{1}$ is constant, its Fourier transform ${\mathscr{Q}}_{1}$ is strictly local and $\omega_{1}=\omega_{{\mathscr{Q}}_{1}}$ is a product state. We conclude that $\omega_{{\mathscr{P}}}\,\hat{\otimes}\,\omega_{\mathrm{empty}}$ is LGA equivalent to a product state. Since $\omega_{\mathrm{empty}}$ is also a product state, we conclude that $\omega_{{\mathscr{P}}}$ is stably SRE, see Section 2.4. ∎
Appendix C Flux insertion for free fermions
We take up the setting of Section 5 and prove the various bounds on free fermion operators on ${\mathscr{K}}_{m}$ used in the arguments presented there. We prove in particular the crucial Proposition C.6 relating the quasi-adiabatic flow and the spectral flow whenever the gap does not close. Recall that we fix an orthonormal basis $\{e_{i}\}_{i=1,\cdots,2m}$ of ${\mathbb{C}}^{2m}$ so that $\{\delta_{x}\otimes e_{i}\}_{x,i}$ is an orthonormal basis of ${\mathcal{K}}_{m}$. For any operator ${\mathscr{A}}$ on ${\mathcal{K}}_{m}$ we denote by
${\mathscr{A}}(x,i;y,j):=\langle\delta_{x}\otimes e_{i},{\mathscr{A}}\,\delta_{y}\otimes e_{j}\rangle$ (C.1)
@@ Appendix C Flux insertion for free fermions
Throughout this appendix we will use the following notations; $L=\{x\in{\mathbb{Z}}^{2}\,:\,x_{1}\leq 0\}$ is the left half plane and $R={\mathbb{Z}}^{2}\setminus L$ is the right half plane. We let $l_{L}=(0,-(1,0))$ be the left horizontal axis and $l_{R}=(0,(1,0))$ the right horizontal axis, regarded as a subsets of ${\mathbb{R}}^{2}$, and $l=l_{L}\cup l_{R}$ for the horizontal axis. We denote by $h=h_{l_{L}}$ the upper half plane. For any subset $S\subset{\mathbb{Z}}^{2}$ we write $\Pi_{S}$ for the projection on S, i.e. $\Pi_{S}|\delta_{x}\otimes i\rangle=\delta_{x\in S}|\delta_{x}\otimes i\rangle$. For any $S\subset{\mathbb{R}}^{2}$ and any $x\in{\mathbb{Z}}^{2}$ we write $\mathrm{dist}(x,S)=\inf_{y\in S}\|x-y\|_{1}$ for the distance from $x$ to the set $S$.
We consider a projector ${\mathscr{P}}$ that is exponentially local, i.e. for all $x,y\in{\mathbb{Z}}^{2}$ and all $i,j$ we have
$\left|{\mathscr{P}}(x,i;y,j)\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}$ (C.2)
for some constants $C,\eta>0$, and such that ${\mathscr{T}}{\mathscr{P}}{\mathscr{T}}^{*}={\mathscr{P}}$. We let ${\mathscr{H}}=\operatorname{\mathbbm{1}}-{\mathscr{P}}$ be the Hamiltonian with gap $\Delta=(0,1)$ and Fermi projector ${\mathscr{P}}=\chi_{(-\infty,\mu]}({\mathscr{H}})$ for any $\mu\in\Delta$. We denote by ${\mathscr{H}}_{\phi}$ the flux Hamiltonians defined in (2.17). We have $\left|{\mathscr{H}}_{\phi}(x,i;y,j)\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}$ for all $\phi$ and all $x,y,i$ and $j$. Recall also that for any $\mu\in\Delta$ we denote by ${\mathscr{P}}^{(\mu)}_{\phi}=\chi_{(-\infty,\mu]}({\mathscr{H}}_{\phi})$ the Fermi projections of ${\mathscr{H}}_{\phi}$ for Fermi energy $\mu$.
Corresponding to the projectors ${\mathscr{P}}^{h}_{\phi}={\mathrm{e}}^{{\mathrm{i}}\phi\Pi_{h}}\,{\mathscr{P}}\,{\mathrm{e}}^{-{\mathrm{i}}\phi\Pi_{h}}$, where $h$ is the upper half-plane, we have Hamiltonians ${\mathscr{H}}_{\phi}^{h}=\operatorname{\mathbbm{1}}-{\mathscr{P}}^{h}_{\phi}={\mathrm{e}}^{{\mathrm{i}}\phi\Pi_{h}}\,{\mathscr{H}}_{\phi}^{h}\,{\mathrm{e}}^{-{\mathrm{i}}\phi\Pi_{h}}$ all with spectral gap $\Delta$.
C.1 Basic locality estimates
Lemma C.1.
@@ C.1 Basic locality estimates
$\left|{\mathscr{P}}_{\phi}^{h}(x,i;y,j)\right|$ $\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}$ (C.3)
$\left|\partial_{\phi}{\mathscr{P}}_{\phi}^{h}(x,i;y,j)\right|$ $\leq c{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}$ (C.4)
$\left|{\mathscr{K}}_{\phi}^{h}(x,i;y,j)\right|$ $\leq c{\mathrm{e}}^{-\xi(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}$ (C.5)
for some constants $c,\xi>0$ and for all $\phi\in{\mathbb{R}}$, all $x=(x_{1},x_{2}),y=(y_{1},y_{2})\in{\mathbb{Z}}^{2}$ and all $i,j\in\{1,\ldots 2m\}$.
@@
The first claim follows immediately from Eq (C.2) and the definition ${\mathscr{P}}_{\phi}^{h}={\mathrm{e}}^{{\mathrm{i}}\phi\Pi_{h}}{\mathscr{P}}{\mathrm{e}}^{-{\mathrm{i}}\phi\Pi_{h}}$.
Let us now bound the matrix elements of $\partial_{\phi}{\mathscr{P}}_{\phi}^{h}={\mathrm{i}}[\Pi_{h},{\mathscr{P}}_{\phi}^{h}]$. If $x,y\in h$ or $x,y\notin h$ then $(\partial_{\phi}{\mathscr{P}}_{\phi}^{h})(x,i;y,j)=0$. If $x\in h$ and $y\not\in h$ then using the first claim of this lemma we obtain
$\left|(\partial_{\phi}{\mathscr{P}}_{\phi}^{h})(x,i;y,j)\right|=\left|{\mathscr{P}}_{\phi}(x,i;y,j)\right|\leq C{\mathrm{e}}^{-\eta\|x-y\|_{1}}\leq C{\mathrm{e}}^{-\eta(\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|)}.$ (C.6)
where for the last inequality we used $x\in h$ and $y\not\in h$ so $\|x-y\|_{1}=\left|x_{1}-y_{1}\right|+\left|x_{2}-y_{2}\right|=\left|x_{1}-y_{1}\right|+\left|x_{2}\right|+\left|y_{2}\right|$. The same result is obtained if $x\not\in h$ and $y\in h$, thus establishing the second bound. The last bound of the lemma now follows straightforwardly from the definition ${\mathscr{K}}_{\phi}^{h}={\mathrm{i}}[\partial_{\phi}{\mathscr{P}}_{\phi}^{h},{\mathscr{P}}_{\phi}^{h}]$ and the bounds on the matrix elements of $\partial_{\phi}{\mathscr{P}}_{\phi}^{h}$ and ${\mathscr{P}}_{\phi}^{h}$. ∎
C.2 Comparison of quasi-adiabatic and spectral generators
Fix a Fermi energy $\mu\in\Delta$ and an interval $[\phi^{(i)},\phi^{(f)}]$ such that $\mu$ is in a spectral gap $\Delta^{\prime}=[\mu-\delta,\mu+\delta]$ of ${\mathscr{H}}_{\phi}$ for all $\phi\in[\phi^{(i)},\phi^{(f)}]$. Then the Kato generators
${\mathscr{G}}^{(\mu)}_{\phi}:=-{\mathrm{i}}[\partial_{\phi}{\mathscr{P}}^{(\mu)}_{\phi},{\mathscr{P}}_{\phi}^{(\mu)}]$ (C.7)
are well-defined for all $\phi\in[\phi^{(i)},\phi^{(f)}]$. We want to show that these generators are close to the generators ${\mathscr{K}}_{\phi}=\Pi_{L}{\mathscr{K}}_{\phi}^{h}\Pi_{L}$. First of all, we show some locality preperties.
Lemma C.2.
There are constants $c,\eta>0$ which depend only on the interval $[\phi^{(i)},\phi^{(f)}]$ such that
$\left|{\mathscr{P}}_{\phi}^{(\mu)}(x,i;y,j)\right|$ $\leq\,c{\mathrm{e}}^{-\xi(\|x-y\|_{1})}$ (C.8)
$\left|\partial_{\phi}{\mathscr{P}}_{\phi}^{(\mu)}(x,i;y,j)\right|$ $\leq\,c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{L}))}$ (C.9)
$\left|({\mathscr{P}}^{(\mu)}_{\phi}-{\mathscr{P}}^{h}_{\phi})(x,i;y,j)\right|,\left|\partial_{\phi}({\mathscr{P}}_{\phi}^{(\mu)}-{\mathscr{P}}^{h}_{\phi})(x,i;y,j)\right|$ $\leq\,c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{R}))}$ (C.10)
$\left|{\mathscr{G}}_{\phi}^{(\mu)}(x,i;y,j)\right|$ $\leq\,c{\mathrm{e}}^{-\xi(\|x-y\|_{1}+\mathrm{dist}(x,l_{L}))}$ (C.11)
for all $x,i,y,j$ and $\phi\in[\phi_{1},\phi_{2}]$.